# Spatio-Temporal Fuzzy-oriented Multi-Modal Meta-Learning for Fine-grained Emotion Recognition

### v2

@@ Abstract
Fine-grained emotion recognition (FER) plays a vital role in various fields, such as disease diagnosis, personalized recommendations, and multimedia mining. However, existing FER methods face three key challenges in real-world applications: (i) they rely on large amounts of continuously annotated data to ensure accuracy since emotions are complex and ambiguous in reality, which is costly and time-consuming; (ii) they cannot capture the temporal heterogeneity caused by changing emotion patterns, because they usually assume that the temporal correlation within sampling periods is the same; (iii) they do not consider the spatial heterogeneity of different FER scenarios, that is, the distribution of emotion information in different data may have bias or interference. To address these challenges, we propose a Spatio-Temporal Fuzzy-oriented Multi-modal Meta-learning framework (ST-F2M). Specifically, ST-F2M first divides the multi-modal videos into multiple views, and each view corresponds to one modality of one emotion. Multiple randomly selected views for the same emotion form a meta-training task. Next, ST-F2M uses an integrated module with spatial and temporal convolutions to encode the data of each task, reflecting the spatial and temporal heterogeneity. Then it adds fuzzy semantic information to each task based on generalized fuzzy rules, which helps handle the complexity and ambiguity of emotions. Finally, ST-F2M learns emotion-related general meta-knowledge through meta-recurrent neural networks to achieve fast and robust fine-grained emotion recognition. Extensive experiments show that ST-F2M outperforms various state-of-the-art methods in terms of accuracy and model efficiency. In addition, we construct ablation studies and further analysis to explore why ST-F2M performs well. The code is publicly available at https://github.com/WangJingyao07/ST-F2M.
@@ I Introduction
Fine-grained emotion recognition (FER) uses physiological signals to model the changes in emotions and provide refined descriptions. FER has various applications in multiple fields such as disease diagnosis [1, 2], social media analysis [3, 4, 5], and market research [6]. Take the application of FER in disease diagnosis as an example, it can assist in diagnosing depression, schizophrenia, and other psychological diseases, and also adjust the medical plan and service based on the patient’s emotional feedback [7, 1, 8].
@@
Figure 1: Illustration of our motivation, i.e., the three key challenges in FER.
Recently, many FER methods have been proposed to simulate the dynamics of human emotional states [9, 10, 11, 12]. Unlike discrete emotion recognition (DER) [13, 14], which only assigns a single discrete emotion label to a static image, FER can capture the nature of human emotions changing over time by using fine-grained segments (usually 0.5 seconds to 4 seconds) [15, 16]. Therefore, FER can achieve higher prediction accuracy than DER [17]. Although significant efforts have been made to improve FER, existing methods still face three key challenges, especially in real-life long videos.
@@ II-A Fine-grained Emotion Recognition
Fine-grained emotion recognition is a key task with many practical applications. It has evolved from physics-based models [26, 27] to data-driven models [28, 29], and from coarse-grained [22, 30] to fine-grained analysis [7]. Different types of data can be used to detect fine-grained emotions, such as body movements [31, 32], blood pressure [33], textual information [34] and facial expressions [35, 36]. Among them, facial expressions and textual information are the most convenient and important sources for emotion recognition [28], as they are easily captured and abundant in social media.
@@ II-C Meta-Learning for Emotion Recognition
Humans can learn from a few samples and easily develop an understanding of new tasks, thus interacting with the external world. For example, humans can quickly identify different types of apples in real life after only seeing a picture of an apple. Meta-learning aims at equipping machines with this ability to adapt to new scenarios effectively [54, 55, 56]. Meta-learning methods can be categorized into three types: optimization-based methods [57, 58, 59], model-based methods [60, 61], and metric-based methods [62, 63, 64]. They all rely on shared structures to extract meta-knowledge, enabling excellent performance on new tasks and are widely applied in few-shot learning.
@@
III Preliminaries
In this section, we first introduce the foundations of meta-learning. Next, we present the details of the proposed generalized fuzzy-based rules for fine-grained emotion recognition and their intensity quantification.
III-A Meta-Learning
@@
Meta-learning can be conceptualized as a two-level optimization problem. In the first level (inner loop), we learn the task-specific model $f^{i}_{\theta}$ for task $\tau_{i}$ with the support set $\mathcal{D}^{s}{i}$ utilizing the meta-learning model $f_{\theta}$. The objective can be presented as:
@@
where $\alpha$ is the learning rate. In the second level (outer loop), we update the meta-learning model $f_{\theta}$ by minimizing the average loss computed across multiple tasks, incorporating the query sets $\mathcal{D}^{q}$ and leveraging the knowledge gained from the previously learned task-specific models. The objective can be structured as:
@@
where $\beta$ is the learning rate. Note that $f_{\theta}^{i}$ is derived by computing the derivative of $f_{\theta}$, making $\nabla_{f_{\theta}}\frac{1}{N_{tr}}\sum_{i=1}^{N_{tr}}\mathcal{L}(X_{i}^{q},Y_{i}^{q},f_{\theta}^{i})$ be interpreted as the second-order derivative of $f_{\theta}$. In this study, we will mine meta-knowledge in fine-grained emotion recognition based on this paradigm to improve generalization and transferability.
@@
III-B Generalized Fuzzy-based Rules
@@
FKIS aims to obtain sequences describing emotional intensity. The fuzzy inference rules within FKIS are constructed based on psychological knowledge [69, 70], and stored in a knowledge-based engine. The engine takes the changes in a set of physiological attributes as input and computes fuzzy sets based on the membership functions for each attribute. Subsequently, the engine generates 18 classification values for (composite) emotional mappings by combining trigger rules and performing de-fuzzification based on the fuzzy sets. Each basic emotion has three intensity levels, i.e., low, medium, and high. Table II provides a set of FKIS rules for determining emotional intensity, with each element representing the current state of a physiological attribute. Figure 3 illustrates the membership functions for the intensity value of six base emotions. Take Figure 3 (a) as an example, when ”Eccentricity=0.2”, the membership degree of “Angry-Low” is 0, and the membership degree of “Angry-Medium” is 0.34. The coding of the emotions refers to representing emotions as numerical or symbolic values, while membership functions are used to describe the degree of belonging of each data point to each fuzzy set, i.e., the extent of each data point in each emotion category, or its degree in each facial component. We weight the emotional intensity encoding (shown in Table II) based on the membership value and introduce it into the model calculation as fuzzy information. The fuzzy membership functions calculate the coding of the emotions for each view and introduce one-dimensional information with fuzzy semantics, which better captures the complexity of emotions. Note that different emotions have varying numbers of rules; for instance, anger has 42 rules, while disgust only has 34 rules.
@@
The introduction of fuzzy rules can help deal with the complexity and ambiguity of emotions in reality, improve accuracy, and speed up reasoning. First, both building a fuzzy inference system and data annotation for the first time require human knowledge as a prior. However, the systems we built are based on fuzzy emotional reasoning rules constructed from human characteristic components, which are universal and applicable to all emotions once and for all. For new data, the features can still be extracted by the feature extractor and then input into the inference system for fuzzy inference without the need to reconstruct the rules. In contrast, data annotation is to identify and annotate videos from the pixel level. New video data needs to be re-annotated, which will incur new costs. In addition, our fuzzy system contains a built-in inference function. We only need to write down the encoding ${-1,0,1}$ of the characteristic components required for the emotion, and then we can perform inference based on the built-in functions in FCIS and FKIS without requiring a separate model training process. Thus, we can generate new fuzzy rules based on the inference function at any time without additional cost. Therefore, the introduction of fuzzy rules greatly improves the model effect and speeds up reasoning, while the cost of first construction is negligible for the model.
In this study, the above systems, i.e., FCIS and FKIS, are used in two ways: (i) for the decision-making process of ST-F2M: followed by the spatio-temporal integrated module to assign fuzzy semantic information containing emotional intensity to the representation; (ii) for dataset construction: assign emotion and intensity information to benchmark datasets in areas such as facial expression recognition to build benchmark datasets for fine-grained emotion recognition. We have constructed two FER datasets, called DISFA-FER and WFLW-FER, based on DISFA [71] and WFLW [72].
IV Methodology
@@
IV-A Meta-learning Task Construction
@@
IV-B Meta-learning Training
@@
IV-B1 Overview of meta-training
@@
Formally, ST-F2M can be defined as $f_{\theta}=h\circ F\circ g$, where $g$ is the shared encoder, $F$ denotes the fuzzy inference, $h$ is the classifier. We first construct the training set $\mathcal{D}_{tr}$ and the test set $\mathcal{D}_{te}$ based on the method in Subsection IV-A, which are obtained from the distribution $p(\mathcal{T})$. For each task $\mathcal{T}_{i}$, we first input it into the encoder $g$ composed of a spatio-temporal convolution integrated module, thereby obtaining the representation $Z_{i}=g(\mathcal{T}_{i})$. Next, $Z_{i}$ is passed into the fuzzy inference system $F$ to obtain the representation $\hat{Z}_{i}=F(Z_{i})$ that introduces fuzzy semantic information. Then, $\hat{Z}_{i}$ will be input into the classification header to generate specific classification results. Ultimately, this process will undergo a bi-level optimization as mentioned in Subsection III-A to learn common meta-knowledge for rapid adaptation to new tasks.
@@ 10       Randomly select $x_{a,b}$ with the same emotional label, obtaining task $\mathcal{T}_{i}$;
17             Perform generalized fuzzy inference with FCIS and FKIS mentioned in Subsection III-B, obtaining $\hat{Z}_{i}=F(Z_{i})$;
@@
IV-B2 Spatio-temporal encoding
@@
IV-B3 Generalized fuzzy inference
For generalized fuzzy inference, we perform secondary encoding based on the fuzzy systems described in Subsection III-B, that is, $\hat{Z}_{i}=F(Z_{i})$.
@@ IV-B4 Overall optimization
Specifically, we perform bi-level optimization as described in Subsection III-A. The objective of the first level optimization (inner loop) is updated to:
@@
V Experiments
In this section, we first introduce the experimental setup in Subsection V-A. Then, we present the results of performance comparison, computational efficiency comparison, and robustness analysis in Subsections V-D to V-C, demonstrating the superiority of ST-F2M. Next, we conduct ablation studies and parameter investigation to further explore how ST-F2M works and what makes ST-F2M work well in Subsections V-E to V-F respectively. Finally, we evaluate the practical application effect of ST-F2M in Subsection V-G.
V-A Experimental Setup
V-A1 Implementation Details
We opt for the most commonly used Conv4 backbone [57] as the foundational structure for the encoder. After the convolution and filtering stages, we sequentially apply batch normalization, ReLU non-linear activation, and $2\times 2$ max pooling (achieved through stride convolutions). The last layer ultimately produces two separate channels as its final output. The first channel is fed into the proposed fuzzy systems, i.e., FCIS and FKIS, to extract fuzzy semantic information, while the second channel is directed towards a softmax layer. These network architectures undergo an initial pre-training phase and remain unaltered throughout the training process. As we transition to the optimization stage, we utilize the Adam optimizer [77] for training our model. We set the values of momentum and weight decay to 0.9 and $10^{-4}$, respectively. The initial learning rate for all experiments is set at 0.2, with the flexibility for linear scaling as needed. All experiments are executed on 10 NVIDIA V100 GPUs.
V-A2 Datasets
V-A3 Baselines
We introduce three types of baselines: (i) state-of-the-art methods in each dataset, including FN2EN [84], VGG-F [85], PropNet-CNN [86], UniMSE [19], and SepTr [87]; (ii) prominent multi-modal FER methods, including PATHOSnet-v2 [88], COGMEN [89], and MARLIN [90]; and (iii) meta-learning methods, including MAML [57], Reptile [91], ProtoNet [92], RelationNet [93], MetaCRL [94], and ANIL [59]. Note that for single-modal methods under multi-modal data sets, we use the bagging [95] method to integrate after solving the single modality. For the meta-learning method, we adopt the same task construction method as ST-F2M for multi-modal emotion recognition.
@@ V-A3 Baselines
V-B Performance Comparison
To evaluate the effectiveness of ST-F2M, we conduct comparative experiments with the baselines on the five datasets mentioned in Section V-A, containing six basic emotions, i.e., happiness, anger, sadness, disgust, fear, and surprise, and each emotion consists of three intensity levels, i.e., low, medium and high. In this experiment, we set two metrics, including accuracy (ACC) and recall (RECALL). We record the average value of 10 rounds to obtain the results of these two metrics.
The results are shown in Table III. Across all datasets, our approach consistently outperforms both meta-learning methods and multi-modal facial expression recognition methods. Moreover, our ST-F2M achieves results similar to state-of-the-art (SOTA) baselines which rely on large datasets, while our ST-F2M only has limited data support. This indicates that ST-F2M can achieve similar or even superior generalization improvements compared to baselines, without the need for extensive data reliance. It is worth noting that our method is always the SOTA method on the Recall, which illustrates its superiority in fields such as disease diagnosis that require high recall rates [96].
@@ V-B Performance Comparison
V-C Computational Efficiency Comparison
V-D Robustness Analysis
Emotion recognition in real life can encounter various interferences. Take medical environments as an example, the emotional data of the patients may have large portions obscured due to wearing masks or bandages. A robust method should still achieve reliable emotion recognition in the presence of various interferences. To assess the robustness of ST-F2M, we manually introduce multiple noises into the data to simulate real-world interferences, including adding fog [97], masks [98], and distortion [99] for visual modality and adding masks for textual modality. All types of noise are introduced into the data with probabilities of 10%, 30%, and 50%, respectively.
@@ V-D Robustness Analysis
V-E Ablation Study

### v4

@@ Abstract
Fine-grained emotion recognition (FER) plays a vital role in various fields, such as disease diagnosis, personalized recommendations, and multimedia mining. However, existing FER methods face three key challenges in real-world applications: (i) they rely on large amounts of continuously annotated data to ensure accuracy since emotions are complex and ambiguous in reality, which is costly and time-consuming; (ii) they cannot capture the temporal heterogeneity caused by changing emotion patterns, because they usually assume that the temporal correlation within sampling periods is the same; (iii) they do not consider the spatial heterogeneity of different FER scenarios, that is, the distribution of emotion information in different data may have bias or interference. To address these challenges, we propose a Spatio-Temporal Fuzzy-oriented Multi-modal Meta-learning framework (ST-F2M). Specifically, ST-F2M first divides the multi-modal videos into multiple views, and each view corresponds to one modality of one emotion. Multiple randomly selected views for the same emotion form a meta-training task. Next, ST-F2M uses an integrated module with spatial and temporal convolutions to encode the data of each task, reflecting the spatial and temporal heterogeneity. Then it adds fuzzy semantic information to each task based on generalized fuzzy rules, which helps handle the complexity and ambiguity of emotions. Finally, ST-F2M learns emotion-related general meta-knowledge through meta-recurrent neural networks to achieve fast and robust fine-grained emotion recognition. Extensive experiments show that ST-F2M outperforms various state-of-the-art methods in terms of accuracy and model efficiency. In addition, we construct ablation studies and further analysis to explore why ST-F2M performs well.
@@
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
Fine-grained emotion recognition (FER) uses physiological signals to model the changes in emotions and provide refined descriptions. FER has various applications in multiple fields such as disease diagnosis [1, 2], social media analysis [3, 4], and market research [5]. Take the application of FER in disease diagnosis as an example, it can assist in diagnosing depression, schizophrenia, and other psychological diseases, and also adjust the medical plan and service based on the patient’s emotional feedback [6, 7, 1, 8].
@@
Figure 1: Illustration of our motivation, i.e., the three key challenges in fine-grained emotion recognition.
Recently, many FER methods have been proposed to simulate the dynamics of human emotional states [9, 10, 11]. Unlike discrete emotion recognition (DER) [12, 13] which only assigns a single discrete emotion label to an image, FER can capture the nature of human emotions changing over time by using fine-grained segments (usually 0.5 seconds to 4 seconds) [14, 15]. Therefore, FER can achieve higher prediction accuracy than DER [16]. Although significant efforts have been made to improve FER performance, existing methods still face three key challenges, especially in real-life long videos.
@@ II-A Fine-grained Emotion Recognition
Fine-grained emotion recognition is a key task with many practical applications. It has evolved from physics-based models [26, 27] to data-driven models [28, 29], and from coarse-grained [21, 30] to fine-grained analysis [22]. Different types of data can be used to detect fine-grained emotions, such as body movements [31], blood pressure [32], textual information [33] and facial expressions [34]. Among them, facial expressions and textual information are the most convenient and important sources for emotion recognition [28], as they are easily captured and abundant in social media.
@@ II-C Meta-Learning for Emotion Recognition
Humans can learn from a few samples and easily develop an understanding of new tasks, thus interacting with the external world. For example, humans can quickly identify different types of apples in real life after only seeing a picture of an apple. Meta-learning aims at equipping machines with this ability to adapt to new scenarios effectively [52]. Meta-learning methods can be categorized into three types: optimization-based methods [53, 54, 55], model-based methods [56, 57], and metric-based methods [58, 59, 60]. They all rely on shared structures to extract meta-knowledge, enabling excellent performance on new tasks, and are widely applied in few-shot learning.
@@
III Problem Definition
In this section, we first introduce the problem settings and notations of meta-learning for FER.
@@ III Problem Definition
Meta-learning can be conceptualized as a two-level optimization problem. In the first level (inner loop), we learn the task-specific model $f^{i}_{\theta}$ for task $\tau_{i}$ with the support set $\mathcal{D}^{s}{i}$ utilizing the meta-learning model $f_{\theta}$. The objective is:
@@
where $\alpha$ is the learning rate. In the second level (outer loop), we update the meta-learning model $f_{\theta}$ by minimizing the average loss computed across multiple tasks, incorporating the query sets $\mathcal{D}^{q}$ and leveraging the knowledge gained from the previously learned task-specific models. The objective is:
@@
where $\beta$ is the learning rate. Note that $f_{\theta}^{i}$ is derived by computing the derivative of $f_{\theta}$, making $\nabla_{f_{\theta}}\frac{1}{N_{tr}}\sum_{i=1}^{N_{tr}}\mathcal{L}(X_{i}^{q},Y_{i}^{q},f_{\theta}^{i})$ be interpreted as the second-order derivative of $f_{\theta}$.
In this study, we will mine meta-knowledge in fine-grained emotion recognition based on this paradigm to improve generalization and transferability.
@@
IV Generalized Fuzzy-based Rules
@@
FKIS aims to obtain sequences describing emotional intensity. The fuzzy inference rules within FKIS are constructed based on psychological knowledge [65, 66], and stored in a knowledge-based engine. The engine takes the changes in a set of physiological attributes as input and computes fuzzy sets based on the membership functions for each attribute. Subsequently, the engine generates 18 classification values for (composite) emotional mappings by combining trigger rules and performing de-fuzzification based on the fuzzy sets. Each basic emotion has three intensity levels, i.e., low, medium, and high. Table II provides a set of FKIS rules for determining emotional intensity, with each element representing the current state of a physiological attribute. Figure 3 illustrates the membership functions for the intensity value of six base emotions. Take Figure 3 (a) as an example, when “Eccentricity=0.2”, the membership degree of “Angry-Low” is 0, and the membership degree of “Angry-Medium” is 0.34. The coding of the emotions refers to representing emotions as numerical or symbolic values, while membership functions are used to describe the degree of belonging of each data point to each fuzzy set, i.e., the extent of each data point in each emotion category, or its degree in each facial component. We weight the emotional intensity encoding (shown in Table II) based on the membership value and introduce it into the model calculation as fuzzy information. The fuzzy membership functions calculate the coding of the emotions for each view and introduce one-dimensional information with fuzzy semantics, which better captures the complexity of emotions. Note that different emotions have varying numbers of rules; for instance, anger has 42 rules, while disgust only has 34 rules.
@@
In this study, the above systems, i.e., FCIS and FKIS, are used in two goals: (i) for the decision-making process of ST-F2M: followed by the spatio-temporal integrated module to assign fuzzy semantic information containing emotional intensity to the representation; (ii) for dataset construction: assign emotion and intensity information to benchmark datasets in areas such as facial expression recognition to build benchmark datasets for fine-grained emotion recognition.
For the first objective, the introduction of fuzzy rules helps deal with the complexity and ambiguity of emotions in reality and improve model generalization. Specifically, the proposed emotional reasoning system is grounded in psychological theories of human affect, wherein complex emotional states are decomposed into a set of elementary characteristic components. These components serve as the minimal semantic units of emotion, analogous to “building blocks”, and enable a compositional representation of emotional expressions. This formulation allows emotional data, regardless of context or modality, to be interpreted as permutations of these basic units, thus supporting generalizability and reusability across different scenarios. Further, the fuzzy membership mechanism allows each component to be activated with varying intensities depending on context and subject, addressing the intrinsic individual variability and ambiguity in emotional expression. This enables the system to model the uncertainty and continuity inherent in human emotions more robustly than discrete categorical methods. The specific implementation is illustrated in Section V.
For the second objective, our system significantly reduces data maintenance overhead compared with traditional annotation-based approaches, which demand costly manual labeling for every new video or sample. First, both building a fuzzy inference system and data annotation for the first time require human knowledge as a prior. However, the systems we built are based on fuzzy emotional reasoning rules constructed from human characteristic components, which are universal and applicable to all emotions once and for all. For new data, the features can still be extracted by the feature extractor and then input into the inference system for fuzzy inference without the need to reconstruct the rules. In contrast, data annotation is to identify and annotate videos from the pixel level. New video data needs to be re-annotated, which will incur new costs. In addition, our fuzzy system contains a built-in inference function. We only need to write down the encoding ${-1,0,1}$ of the characteristic components required for the emotion, and then we can perform inference based on the built-in functions in FCIS and FKIS without requiring a separate model training process. Thus, we can generate new fuzzy rules based on the inference function at any time without additional cost. Therefore, the introduction of fuzzy rules greatly improves the model effect and speeds up reasoning, while the cost of first construction is negligible for the model. We have constructed two FER datasets, called DISFA-FER and WFLW-FER, based on DISFA [67] and WFLW [68].
V Methodology
@@
V-A Meta-learning Task Construction
@@
V-B Meta-learning Training
@@
V-B1 Overview of meta-training
@@
Formally, ST-F2M can be defined as $f_{\theta}=h\circ F\circ g$, where $g$ is the shared encoder, $F$ denotes the fuzzy inference, $h$ is the classifier. We first construct the training set $\mathcal{D}_{tr}$ and the test set $\mathcal{D}_{te}$ based on the method in Subsection V-A, which are obtained from the distribution $p(\mathcal{T})$. For each task $\mathcal{T}_{i}$, we first input it into the encoder $g$ composed of a spatio-temporal convolution integrated module, thereby obtaining the representation $Z_{i}=g(\mathcal{T}_{i})$. Next, $Z_{i}$ is passed into the fuzzy inference system $F$ to obtain the representation $\hat{Z}_{i}=F(Z_{i})$ that introduces fuzzy semantic information. Then, $\hat{Z}_{i}$ will be input into the classification header to generate specific classification results. Ultimately, this process will undergo a bi-level optimization as mentioned in Subsection III to learn common meta-knowledge for rapid adaptation to new tasks.
@@ 10       Randomly select $x_{a,b}$ with the same emotional label, obtaining task $\mathcal{T}_{i}$;
17             Perform generalized fuzzy inference with FCIS and FKIS mentioned in Subsection IV, obtaining $\hat{Z}_{i}=F(Z_{i})$;
@@
V-B2 Spatio-temporal encoding
@@
V-B3 Generalized fuzzy inference
For generalized fuzzy inference, we perform secondary encoding based on the fuzzy systems described in Subsection IV, that is, $\hat{Z}_{i}=F(Z_{i})$.
@@ V-B4 Overall optimization
Specifically, we perform bi-level optimization as described in Subsection III. The objective of the first level optimization (inner loop) is updated to:
@@
TABLE V: Performance comparison when facing mislabeling with few-shot settings on CK+. “Standard” indicates the general settings as in Subsection VI-B, and “Few-Shot” indicates the few-shot settings with 30% mislabeled samples.
Methods Standard Few-Shot
Average ACC Worst ACC Average ACC Worst ACC
MAML 85.70 $\pm$ 0.64 83.16 $\pm$ 0.59 49.14 $\pm$ 0.56 45.12 $\pm$ 0.77
ProtoNet 86.40 $\pm$ 0.67 84.21 $\pm$ 0.58 48.15 $\pm$ 0.62 44.37 $\pm$ 0.60
PATHOSnet-v2 92.74 $\pm$ 0.64 91.30 $\pm$ 0.44 43.15 $\pm$ 0.27 40.50 $\pm$ 0.36
MARLIN 94.76 $\pm$ 0.14 92.8 $\pm$ 0.36 45.41 $\pm$ 0.22 42.01 $\pm$ 0.63
BMHP 94.15 $\pm$ 0.19 90.69 $\pm$ 0.25 48.58 $\pm$ 0.32 42.79 $\pm$ 0.45
FPM-Net 95.02 $\pm$ 0.56 93.46 $\pm$ 0.54 49.28 $\pm$ 0.32 47.28 $\pm$ 0.38
TAE 95.19 $\pm$ 0.77 91.49 $\pm$ 0.84 50.41 $\pm$ 0.65 47.72 $\pm$ 0.58
CBERL 94.30 $\pm$ 0.25 91.76 $\pm$ 0.18 51.41 $\pm$ 0.83 48.22 $\pm$ 0.78
Ours 96.96 $\pm$ 0.52 95.01 $\pm$ 0.39 57.46 $\pm$ 0.50 55.80 $\pm$ 0.47
@@
VI Experiments
In this section, we first introduce the experimental setup in Subsection VI-A. Then, we present the results of performance comparison, computational efficiency comparison, and robustness analysis in Subsections VI-D to VI-C, demonstrating the superiority of ST-F2M. Next, we conduct ablation studies and parameter investigation to further explore how ST-F2M works and what makes ST-F2M work well in Subsections VI-E to VI-F respectively. Finally, we evaluate the practical application effect of ST-F2M in Subsection VI-G.
VI-A Experimental Setup
VI-A1 Implementation Details
We opt for the most commonly used ResNet-18 as the backbone. After the convolution and filtering stages, we sequentially apply batch normalization, ReLU non-linear activation, and $2\times 2$ max pooling (achieved through stride convolutions). The last layer ultimately produces two separate channels as its final output. The first channel is fed into the proposed fuzzy systems, i.e., FCIS and FKIS, to extract fuzzy semantic information, while the second channel is directed towards a softmax layer. These network architectures undergo an initial pre-training phase and remain unaltered throughout the training process. As we transition to the optimization stage, we utilize the Adam optimizer [73] for training our model. We set the values of momentum and weight decay to 0.9 and $10^{-4}$, respectively. The initial learning rate for all experiments is set at 0.2, with the flexibility for linear scaling as needed. All experiments are executed on 10 NVIDIA V100 GPUs.
VI-A2 Datasets
VI-A3 Baselines
We introduce four types of baselines: (i) state-of-the-art methods (before 2024) of each dataset, including FN2EN [80], VGG-F [81], PropNet-CNN [82], UniMSE [18], and SepTr [83]; (ii) meta-learning methods, including MAML [53], Reptile [84], ProtoNet [85], RelationNet [86], MetaCRL [87], and ANIL [55]; and (iii) prominent and recently proposed multi-modal FER methods, including PATHOSnet-v2 [88], COGMEN [89], and MARLIN [90], BMHP [91], FPM-Net [92], TAE [93], and CBERL [94]. Note that for single-modal methods under multi-modal data sets, we use the bagging [95] method to integrate after solving the single modality. For the meta-learning method, we adopt the same task construction method as ST-F2M for multi-modal emotion recognition.
VI-B Performance Comparison
To evaluate the effectiveness of ST-F2M, we conduct comparative experiments with the baselines on the five datasets mentioned in Section VI-A, containing six basic emotions, i.e., happiness, anger, sadness, disgust, fear, and surprise, and each emotion consists of three intensity levels, i.e., low, medium and high. In this experiment, we set two metrics, including accuracy (ACC) and recall (RECALL). We record the average value of 10 rounds to obtain the results of these two metrics.
The results are shown in Table III. From the results, we can observe that across all datasets, our approach consistently outperforms both meta-learning methods and multi-modal facial expression recognition methods. Moreover, our ST-F2M achieves results similar to state-of-the-art (SOTA) baselines which rely on large datasets, while our ST-F2M only has limited data support. This indicates that ST-F2M can achieve similar or even superior generalization improvements compared to baselines, without the need for extensive data reliance. It is worth noting that our method is always the SOTA method on the Recall, which illustrates its superiority in fields such as disease diagnosis that require high recall rates [96].
VI-C Computational Efficiency Comparison
VI-D Robustness Analysis
VI-D1 When Facing Distortion and Noise
Emotion recognition in real life can encounter various interferences. Take medical environments as an example, the emotional data of the patients may have large portions obscured due to wearing masks or bandages. A robust method should still achieve reliable emotion recognition in the presence of various interferences. To assess the robustness of ST-F2M, we manually introduce multiple noises into the data to simulate real-world interferences, including adding fog [97], masks [98], and distortion [99] for visual modality and adding masks for textual modality. All types of noise are introduced into the data with the levels of 10%, 30%, and 50%, respectively. Specifically, in the case of fogging, these levels also correspond to increasing fog density; in the case of masking, these levels indicate the proportion of the sample area that is masked; in the case of distortion, these levels refer to the proportion of samples that affected.
VI-D2 When Facing Mislabeling with Few Samples
To further assess the robustness of the proposed method, we conduct an experiment targeting scenarios with limited samples and mislabeled data. Using CK+ as the benchmark dataset, we follow the same experimental setup as in Subsection VI-B, except that each training batch contains only 20% of the original data to simulate a few-shot setting. Additionally, 30% of the sampled training data have their emotion labels randomly shuffled to mimic real-world weak supervision with label noise. We report both the average and worst-case test performance of different models in Table V. The results show that our method consistently achieves superior performance under these challenging conditions, demonstrating its effectiveness. This also confirms that the fuzzy system accurately models fundamental emotional components and effectively mitigates the impact of mislabeled data during training.
VI-E Ablation Study